# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ A.2.2 Stationary Regressor
Recall $\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$ as defined in (3.5), and observe the following decomposition
$\zeta_{j,t}=w_{j,t}+\dfrac{C_{\zeta}}{n^{\tau}}\phi_{j,t},\ \phi_{j,t}:=\sum_{s=1}^{t}\rho_{\zeta}^{t-s}w_{j,s-1}.$ (A.43)
@@
In addition, we know $n^{-1}\sum_{t=1}^{n}\phi_{j,t}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}n^{\tau}$ from (A.17). By (A.15), the cross-product between a mildly integrated $\phi_{j,t}$ and a stationary $w_{j,t}$ is bounded by
@@
where the second inequality applies the rate in (A.16) of the cross product between a unit root and a mildly integrated series. Thus,
@@
$\|n^{-1}\sum_{t=1}^{n}X_{-j,t-1}\eta_{j,t}^{(1)}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{1+\frac{1}{2r}}.$
@@
[…]
with $\mathcal{B}_{j}$ being a Brownian motion of variance ${\rm var}\left(e_{j,t}\right)$, and $\eta_{j,t}^{(1)}$ defined in (A.54). In addition,
@@
CASE I: $j\in\mathcal{M}_{x}$. Define $\check{r}_{j,t}=\zeta_{j,t}-W_{-j,t}^{\top}\widehat{\varphi}^{(j)},$ and by definition we have
$\check{r}_{j,t}=\widehat{r}_{j,t}\widehat{\varsigma}_{j}.$ (A.67)
@@
$\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}=\sum_{t=1}^{n}\zeta_{j,t-1}w_{j,t-1}-\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}.$ (A.68)
Note that $w_{j,t}$ is unit root and $\zeta_{j,t}$ is mildly integrated. By the functional CLT for the case of unit root in Phillips and Lee, (2016, Lemma 3.2), we have
[…]
Note that in (A.36), we allow for a small $C_{\varphi}$ that shrinks to zero as $n\to\infty$. Let $C_{\varphi}=1/(4C_{\sigma}s^{2}n^{1-\tau})$ with the absolute constant $C_{\sigma}$ in (A.36). Then
$\left|n^{-(1+\tau)}\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}\right|$ $\leq\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\|D_{-j}^{-1}\widehat{\varphi}^{(j)}\|_{1}$
$\leq\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\dfrac{4C_{\sigma}C_{\varphi}s^{2}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\leq\dfrac{\sup_{k\in[p],t\in[n]}|w_{k,t-1}|\cdot|w_{j,t-1}|}{n^{\tau}\min_{k\in[p]}\widehat{\sigma}_{k}}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{n^{1-\tau}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\leq\dfrac{n(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{n^{\tau}\cdot n^{1-\tau}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}}=\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\to 0,$
where the second inequality applies (A.36), the fourth inequality applies (A.11), (A.12), (A.9), and (A.39). Thus,
@@
For (A.60), we have $\check{r}_{j,t}^{2}=\zeta_{j,t}^{2}+(W_{-j,t}^{\top}\widehat{\varphi}^{(j)})^{2}-2\zeta_{j,t}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$ When $j\in\mathcal{M}_{x}$,
$\left|\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}-\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\right|$ $\leq\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}^{2}\cdot\|n^{-1}\widetilde{W}_{-j,\cdot}^{\top}\widetilde{W}_{-j,\cdot}\|_{\infty}$
$\ \ \ \ +2\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\cdot\|\dfrac{1}{n}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}\zeta_{j,t}\|_{\infty}.$ (A.72)
@@
$=O\left(n^{\tau-1/2}(\log p)^{\frac{3}{2}+\frac{1}{2r}}\right).$ (A.74)
where the second inequality applies (A.15), (A.16), (A.11), and (A.12). Combining (A.72), (A.73), (A.74), and Proposition 1, we have
$\left|\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}-\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\right|$ $=O_{p}\left(n^{-\tau\wedge(1-\tau)}(\log p)^{4+1/r}+n^{-\frac{\tau\wedge(1-\tau)}{2}+(\tau-1/2)}(\log p)^{\frac{5}{2}+\frac{1}{r}}\right)$
@@
$\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}=\frac{1}{n\widehat{\varsigma}_{j}}\sum_{t=1}^{n}w_{j,t-1}(w_{j,t-1}-Z_{-j,t-1}^{\top}\varphi_{0z}^{(j)*}).$
@@
$\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm cov}\left(w_{j,t},\eta_{j,t}\right)}{\sqrt{{\rm var}\left(w_{j,t}\right)}}$ (A.79)
where $\eta_{j,t}=w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*}$. In addition, we deduce that
@@
where the second step follows (A.10) bounding the cross product between unit roots and a stationary component, (A.11) and (A.12) bounding the standard deviations, and the fact that
@@
$\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm cov}\left(w_{j,t},\eta_{j,t}\right)}{\sqrt{{\rm var}\left(w_{j,t}\right)}}.$ (A.81)
@@
$\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{\left|{\rm cov}\left(w_{j,t},\eta_{j,t}\right)\right|}{{\rm cov}\left(w_{j,t},\eta_{j,t}\right)}={\rm sgn}\left({\rm cov}\left(w_{j,t},\eta_{j,t}\right)\right),$
@@
For (A.60), note that $\widehat{r}_{j,t}$ is the LASSO residual of regression (2.9). Recall that $\widetilde{\eta}_{j,t}$ is the error term of he pseudo-true model (3.10). Following the ideas in the proof of (A.65), we can show that $n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}$ and $n^{-1}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}^{2}$ share the same probability limit, which is
${\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}={\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}^{2}={\rm plim}_{n\to\infty}\dfrac{\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}}{\widehat{\zeta}_{j}^{2}}=\dfrac{{\rm var}\left(\eta_{j,t}\right)}{{\rm var}\left(w_{j,t}\right)},$
where the probability limit of the denominator follows (A.52). This completes the proof of Lemma A.5. ∎
We then prove the main results in Section 3.2
@@
$\mathcal{Z}_{j}$ $=\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\cdot\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\text{$\sigma_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}},$
$\varDelta_{j}$ $=\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})}{\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}.$
@@
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{s^{2}(\log p)^{8+\frac{1}{r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\to 0$
@@
It then suffices to show that $\mathcal{Z}_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ Given the limit established (A.58), when $j\in\mathcal{M}_{z}$ it suffices to show
@@
$\mathcal{Z}_{j}$ $=\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\cdot\mathcal{Z}_{j}^{(1)}$
@@
$\left|\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-1/2}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}u_{t}\|_{\infty}\cdot\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{3/2+1/(2r)}/\sqrt{n^{(1-\tau)\wedge\tau}}\to 0,$
where the first step applies (A.83), the second step applies (A.13) and Proposition 1.
As for the second term, by the central limit theorem in Kostakis et al., (2015, Lemma B4(ii)), the law of large numbers Eq.(21) in the supplement of the same reference, and the Slutsky’s Theorem, we have
@@
Besides, Phillips and Magdalinos, (2009)’s Lemma 3.2 shows that the asymptotic distribution in (A.85) is independent of the Brownian motion in the expression (A.59) of $G_{j}^{*}$. By (A.76) and the Slutsky’s Theorem, we have
$\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\sigma_{u}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}=\sqrt{\dfrac{\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}}\cdot\dfrac{n^{-1/2}\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}}{\sigma_{u}\sqrt{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$

### v3

@@ A.2.2 Stationary Regressor
Recall $\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$ as defined in (3.5), and by (23) in phillips2009econometric we have the following decomposition
$\zeta_{j,t}=w_{j,t}-\dfrac{C_{\zeta}}{n^{\tau}}\phi_{j,t},\ \phi_{j,t}:=\sum_{s=1}^{t}\rho_{\zeta}^{t-s}w_{j,s-1}.$ (A.82)
@@
In addition, following the same arguments for (A.43), we have
$n^{-1}\sum_{t=1}^{n}\phi_{j,t}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}n^{\tau}.$
By (A.41), the cross-product between a mildly integrated $\phi_{j,t}$ and a stationary $w_{j,t}$ is bounded by
@@
where the second inequality applies the rate in (A.42) of the cross product between a local unit root and a mildly integrated series. Thus,
@@
$\|n^{-1}\sum_{t=1}^{n}X_{t-1}\eta_{j,t}^{(1)}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{1+\frac{1}{2r}}.$
@@ A.3.1 Technical Lemmas
$G_{j}^{*}=\begin{cases}\dfrac{1}{C_{\zeta}}\left({\rm lvar}\left(e_{j,t}\right)+\int_{0}^{1}\mathcal{U}_{j}(r)d\mathcal{U}_{j}(r)\right),&j\in\mathcal{M}_{x},\\ {\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right),&j\in\mathcal{M}_{z},\end{cases}$ (A.98)
with ${\rm lvar}\left(e_{j,t}\right)$ being the long-run variance of $e_{j,t}$, $\mathcal{U}_{j}(r)=\int_{0}^{1}{\rm e}^{c_{j}^{*}(r-s)}{\rm d}\mathcal{B}_{j}(s)$ being an OU process, $\mathcal{B}_{j}$ being the Brownian motion of variance ${\rm lvar}\left(e_{j,t}\right)$, and $\eta_{j,t}^{(1)}$ defined in (A.93). In addition,
@@
CASE I: $j\in\mathcal{M}_{x}$. Define
$\check{r}_{j,t}=\widehat{r}_{j,t}\widehat{\varsigma}_{j},$ (A.106)
and by definition of $\widehat{r}_{j,t}$ in (2.11) we have
$\check{r}_{j,t}=\zeta_{j,t}-\widehat{\varsigma}_{j}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}.$ (A.107)
@@
$\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}=\sum_{t=1}^{n}\zeta_{j,t-1}w_{j,t-1}-\widehat{\varsigma}_{j}\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}.$ (A.108)
Note that $w_{j,t}$ is unit root and $\zeta_{j,t}$ is the IV. By the functional CLT for the case of local unit roots in phillips2016robust, we have
$\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}G_{j}^{*}:=\dfrac{1}{C_{\zeta}}\left({\rm lvar}(e_{j,t})+\int_{0}^{1}\mathcal{U}_{j}d\mathcal{U}_{j}\right).$ (A.109)
By the bound of $\min_{k\in[p]}\widehat{\sigma}_{k}$ by (A.35) and (A.36), the bound of LUR processes (A.25), and the bound of a stationary component (A.78), we have
$\sup_{t\in[n]}\|D_{-j}^{-1}W_{-j,t-1}\|_{\infty}\leq\dfrac{\sup_{k\in\mathcal{M}_{x}}|w_{k,t}|}{\min_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}}+\dfrac{\sup_{k\in\mathcal{M}_{z}}|w_{k,t}|}{\min_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}}=O_{p}(\log p).$ (A.110)
Note that in (A.75), we allow for a small $C_{\varphi}$ that shrinks to zero as $n\to\infty$. Let $C_{\varphi}=1/(4C_{\sigma}s^{2}n^{1-\tau/2})$ with the absolute constant $C_{\sigma}$ in (A.75). Then
$\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq\dfrac{4C_{\sigma}C_{\varphi}s^{2}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\leq\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{n^{1-\tau/2}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}},$ (A.111)
where the first inequality applies (A.75). Therefore,
$\left|n^{-(1+\tau)}\widehat{\varsigma}_{j}\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}\right|$ $\leq\widehat{\varsigma}_{j}\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\leq\widehat{\varsigma}_{j}\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\dfrac{4C_{\sigma}C_{\varphi}s^{2}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\leq\dfrac{\widehat{\varsigma}_{j}\sup_{t\in[n]}\|D_{-j}^{-1}W_{-j,t-1}\|_{\infty}|w_{j,t-1}|}{n^{\tau}}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{n^{1-\tau/2}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{n^{\tau/2}\cdot\sqrt{n}(\log p)^{\frac{3}{2}}}{n^{\tau}\cdot}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{n^{1-\tau/2}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}}=\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\to 0,$
where the fourth inequality applies $\widehat{\varsigma}_{j}=O_{p}(n^{\tau/2})$ in (A.43), and the bound of $\sup_{t\in[n]}\|D_{-j}^{-1}W_{-j,t-1}\|_{\infty}$ in (A.110). Thus,
@@
For (A.99), we have $\check{r}_{j,t}^{2}=\zeta_{j,t}^{2}+\widehat{\varsigma}_{j}^{2}(W_{-j,t}^{\top}\widehat{\varphi}^{(j)})^{2}-2\widehat{\varsigma}_{j}\zeta_{j,t}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$ . When $j\in\mathcal{M}_{x}$,
$\left|\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}-\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\right|$ $\leq\widehat{\varsigma}_{j}^{2}\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}^{2}\cdot\|n^{-1}\widetilde{W}_{-j,\cdot}^{\top}\widetilde{W}_{-j,\cdot}\|_{\infty}$
$\ \ \ \ +2\widehat{\varsigma}_{j}\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\cdot\|\dfrac{1}{n}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}\zeta_{j,t}\|_{\infty}.$ (A.114)
@@
$=O\left(n^{\tau-1/2}(\log p)^{\frac{3}{2}+\frac{1}{2r}}\right)=o_{p}(n^{\tau/2}).$ (A.116)
where the second inequality applies (A.41), (A.42), (A.35), and (A.36). Combining (A.114), (A.115), (A.116), the rate of $\widehat{\varsigma}_{j}$ in (A.43), and Proposition 1, we have
$\left|\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}-\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\right|$ $=O_{p}\left(n^{\tau-\tau\wedge(1-\tau)}(\log p)^{4+1/r}\right)+O_{p}(n^{\tau/2})\cdot o_{p}(n^{\tau/2})$
@@
$\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}$ $=\frac{1}{n\widehat{\varsigma}_{j}}\sum_{t=1}^{n}w_{j,t-1}(w_{j,t-1}-Z_{-j,t-1}^{\top}\varphi_{0z}^{(j)*})$
$=\frac{1}{n\widehat{\varsigma}_{j}}\sum_{t=1}^{n}w_{j,t-1}\eta_{j,t-1}^{(1)}.$
@@
$\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)}{\sqrt{{\rm var}\left(w_{j,t}\right)}}$ (A.121)
where $\eta_{j,t}^{(1)}=w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*}$ was defined in (A.93). In addition, we deduce that
@@
where the second step follows (A.38) bounding the cross product between LURs and a stationary component, (A.35) and (A.35) bounding the standard deviations, and the fact that
@@
$\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)}{\sqrt{{\rm var}\left(w_{j,t}\right)}}.$ (A.123)
@@
$\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{\left|{\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)\right|}{{\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)}={\rm sgn}\left({\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)\right),$
@@
For (A.99), note that $\widehat{r}_{j,t}$ is the LASSO residual of regression (2.11). Recall that $\widetilde{\eta}_{j,t}$ is the error term of the pseudo-true model (3.14). Following the ideas in the proof of (A.104), we can show that $n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}$ and $n^{-1}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}^{2}$ share the same probability limit, which is
${\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}$ $={\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}^{2}.$
$={\rm plim}_{n\to\infty}\dfrac{\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}}{\widehat{\varsigma}_{j}^{2}}=\dfrac{{\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}}{{\rm var}(w_{j,t})},$ (A.124)
where the last step applies (A.91). In addition, note that
$\eta_{j,t}$ $=\eta_{j,t-1}^{(1)}+\zeta_{j,t}-w_{j,t}=\eta_{j,t-1}^{(1)}-n^{-\tau}C_{\zeta}\phi_{j,t-1},$ (A.125)
where the first equality applies (A.94), and the second equality applies (A.82). Therefore,
$\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}$ $=\dfrac{1}{n}\sum_{t=1}^{n}(\eta_{j,t-1}^{(1)})^{2}+\dfrac{C_{\zeta}^{2}}{n^{2\tau+1}}\sum_{t=1}^{n}\phi_{j,t-1}^{2}-\dfrac{2C_{\zeta}}{n^{1+\tau}}\sum_{t=1}^{n}\phi_{j,t-1}\eta_{j,t-1}^{(1)}$
$=\dfrac{1}{n}\sum_{t=1}^{n}(\eta_{j,t-1}^{(1)})^{2}+o_{p}(1)$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm var}(\eta_{j,t}^{(1)}),$
where the second row applies the same arguments for (A.89). Then we have
${\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}=\dfrac{{\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}}{{\rm var}(w_{j,t})}=\dfrac{{\rm var}(\eta_{j,t}^{(1)})}{{\rm var}(w_{j,t})}.$
This completes the proof of Lemma A.12. ∎
A.3.2 Proofs of main results in Section 3.3
@@
$\mathcal{Z}_{j}$ $=\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\cdot\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\text{$\sigma_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}},$ (A.126)
$\varDelta_{j}$ $=\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})}{\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}.$ (A.127)
@@
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{s^{2}(\log p)^{8+\frac{1}{r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\to 0$ (A.129)
@@
It then suffices to show that $\mathcal{Z}_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ Given the limit in (A.97), when $j\in\mathcal{M}_{z}$ it suffices to show
@@
$\mathcal{Z}_{j}$ $=\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\cdot\mathcal{Z}_{j}^{(1)}$ (A.132)
@@
$\left|n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-1/2}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}u_{t}\|_{\infty}\cdot\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{3/2+1/(2r)}/\sqrt{n^{(1-\tau)\wedge\tau}},$ (A.133)
where the second step applies (3.8) and Proposition 1. Therefore, by (A.99)
$\dfrac{\left|n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}\right|}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{3/2+1/(2r)}/\sqrt{n^{(1-\tau)\wedge\tau}}\to 0,$
which implies
$\mathcal{Z}_{j}^{(1)}=\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+o_{p}(1)\cdot$ (A.134)
In addition, by the central limit theorem in Lemma B4(ii) of kostakis2015robust and Eq. (28) in their appendix, the law of large numbers Eq. (13) and (21) in the appendix of the same reference, and the Slutsky’s Theorem, we have
@@
Besides, phillips2009econometric’s Lemma 3.2 shows that the asymptotic distribution in (A.135) is independent of the Brownian motion in the expression (A.98) of $G_{j}^{*}$. Also, recall that $\widetilde{\zeta}_{j,t-1}=\zeta_{j,t-1}\widehat{\varsigma}_{j}$ and $\check{r}_{j,t-1}=\widehat{r}_{j,t-1}\widehat{\varsigma}_{j}.$ By (A.118) and the Slutsky’s Theorem, we have
$\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\sigma_{u}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}=\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\sigma_{u}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\cdot\dfrac{\widehat{\varsigma}_{j}}{\widehat{\varsigma}_{j}}=\sqrt{\dfrac{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}}\cdot\dfrac{n^{-1/2}\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}}{\sigma_{u}\sqrt{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$